# Independence Assumption RLS Regularized Least Squares

### v1

@@ 7 Conclusions
Appendix A Related works
@@
A.1 Training models with pairs of responses
Wen et al. (2022), Osband et al. (2023), Johnson et al. (2024) show that we can distinguish between epistemic and aleatoric uncertainty if we train the models with paired observations.
A.2 Epistemic neural nets
@@
A.3 Hallucination detection using first-order methods
@@
A.3.1 Asking language models to quantify uncertainty (self-verification)
@@
A.4 Uncertainty estimation based on sensitivity to contexts
Kassner and Schütze (2020) show that an LLM’s responses can be influenced by irrelevant contexts. Longpre et al. (2021), Neeman et al. (2022) study two sources of knowledge: parametric knowledge stored in the network weights, and contextual knowledge retrieved from external sources. They view reliance of the model on its parametric knowledge and ignoring relevant contextual information as hallucination. These works are mainly motivated by situations where the LLM’s knowledge is outdated and it is instructed to use the (new) contextual information. Accordingly, they design strategies to prioritize contextual information over parametric knowledge. Longpre et al. (2021) also show that larger models are more likely to ignore in-context information in favor of in-weight information. They propose creating training data with modified contextual information so that the model learns to favor the contextual information. Neeman et al. (2022) propose to train a model that predicts two answers: one based on parametric knowledge and one based on contextual information.
@@ A.4 Uncertainty estimation based on sensitivity to contexts
A.5 Hallucination detection using internal states of LLMs
There are a number of papers that try to extract knowledge/truthfulness by inspecting hidden-layer activations of LLMs (Burns et al., 2023, Azaria and Mitchell, 2023, Chen et al., 2024). Such methods clearly require access to the LLM’s internal states, which is not always possible, and severely limits the applicability of these methods.
Appendix B Omitted proofs
Proof of Theorem˜4.5.
@@ Appendix B Omitted proofs
Appendix C Estimation of mutual information and missing mass problem
@@
C.1 The missing mass problem
Let ${\mathcal{X}}$ be a countable set and suppose that $X_{1},\ldots,X_{k}\sim\mu\in{\mathcal{M}}_{1}({\mathcal{X}}^{n})$ independently. In the following $x$ is used as an element of ${\mathcal{X}}^{n}$ rather than the query (as in Section˜4). Then, the missing mass is defined as the random variable
@@ C.1 The missing mass problem
Theorem C.1 (Concentration of a missing mass (Berend and Kontorovich, 2013)).
@@
Note that these estimates are very pessimistic, and in reality we expect the expected missing mass to be significantly smaller. Since natural (and many artificial) languages follow a Zipf distribution (Piantadosi, 2014), we expect that $\mathbb{E}[U_{k}]$ should be much smaller than in the above cases, since sampling from the tail of a Zipf distribution is a rare event. In Section˜C.4 we show the following:
Corollary C.3 (Expected missing mass of Zipf distribution).
@@
The statement followss by combining Lemmas˜C.7 and C.8. ∎
C.2 Estimating mutual information from the partial support
@@
by only having access to $X_{1},\ldots,X_{k}\sim\mu$. Note that that the sample might cover only some part of the support of ${\mathcal{X}}$ and therefore we are facing a missing mass problem. In the following we consider estimator $\widehat{I}_{k}(\gamma)$ given by Algorithm˜1.
In particular in Section˜C.3 we show the following
Theorem C.4.
Fix $\tilde{\mathcal{X}}\subseteq{\mathcal{X}}^{n}$. For any fixed $\gamma>0,\delta\in(0,1)$, with probability at least $1-\delta$,
$(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma)-\left(|\tilde{\mathcal{X}}|\gamma+\ln\left(e+\frac{e}{\gamma}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+\varepsilon_{k}\right)\right)\leq I(\mu)$
@@
In particular, Theorem˜C.4 implies the following:
Corollary C.5.
Under conditions of Theorem˜C.4, there exists $\gamma^{*}_{k}\in(0,1)$ such that
$(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma^{*}_{k})-\left(\frac{1}{k}+(1+n\,\ln\big{(}1+k\,|{\mathcal{X}}|)\big{)}\,\varepsilon_{k}\right)\leq I(\mu)~.$
Note that, choosing any of the upper bounds on $\mathbb{E}U_{k}$ discussed in Section˜C.1, we can see that Corollary˜C.5 implies asymptotic convergence in as a sense
$\lim_{k\to\infty}\widehat{I}_{k}(\gamma^{*}_{k})\leq I(\mu)~.$
C.3 Proof of Theorem˜C.4
@@
Recalling that $S=\big{\{}i\in[k]~:~X_{i}\neq X_{j}\quad\forall j<i\big{\}}$, this immediately implies the following connection between $U_{k}$ and the quantities used in Algorithm˜1:
Proposition C.6.
@@
Note that we use $\sum_{x^{\backslash i}}\mu(\cdots)$ instead of $\mu(x_{i})$ since these are not necessarily equal for some $\mu$.
Introducing a one-dimensional version of $\xi$, for $i\in[n]$ and $z\in{\mathcal{X}}$, as
$\xi_{i}(z)=\mathbb{I}\{z\not\in\{X_{1,i},\ldots,X_{k,i}\}\}~,$
we introduce the normalization factors
$Z=\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\mu(x)~,\qquad Z^{\otimes}=\prod_{i=1}^{n}\sum_{z\in{\mathcal{X}}}(1-\xi_{i}(z))\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},z,x_{i+1},\ldots,x_{n})~.$
We first note that by the definition of $\widehat{\mu}^{\otimes}$, for any $x^{\prime}\in{\mathcal{X}}^{n}$,
$\widehat{\mu}^{\otimes}(x^{\prime})$ $=\prod_{i=1}^{n}\frac{\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},x_{i}^{\prime},x_{i+1},\ldots,x_{n})}{\sum_{j\in S}\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},X_{j,i},x_{i+1},\ldots,x_{n})}$
$=\prod_{i=1}^{n}\frac{\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},x_{i}^{\prime},x_{i+1},\ldots,x_{n})}{\sum_{z\in{\mathcal{X}}}(1-\xi_{i}(z))\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},z,x_{i+1},\ldots,x_{n})}$
$=\frac{\mu^{\otimes}(x^{\prime})}{Z^{\otimes}}~,$
where the second equality comes from the definition of $\xi_{i}$ and the fact that the $X_{i}$ are all different. Now, using the definitions of $\widehat{I}_{k}$ and $\widehat{\mu}$,
$\widehat{I}_{k}(\gamma)$ $=\frac{1}{\sum_{j\in S}\mu(X_{j})}\sum_{i\in S}\mu(X_{i})\left(\ln\left(\frac{\mu(X_{i})}{\sum_{j\in S}\mu(X_{j})}+\frac{\gamma}{\sum_{j\in S}\mu(X_{j})}\right)-\ln\left(\widehat{\mu}^{\otimes}(X_{i})+\frac{\gamma}{Z^{\otimes}}\right)\right)$
$=\frac{1}{Z}\sum_{i\in S}\mu(X_{i})\left(\ln\left(\frac{\mu(X_{i})+\gamma}{Z}\right)-\ln\left(\frac{\mu^{\otimes}(X_{i})+\gamma}{Z^{\otimes}}\right)\right)$ (by Proposition C.6)
$=\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\,\mu(x)\left(\ln\left(\frac{\mu(x)+\gamma}{Z}\right)-\ln\left(\frac{\mu^{\otimes}(x)+\gamma}{Z^{\otimes}}\right)\right)$ (by Eq. 2 )
$\leq\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\,\mu(x)\left(\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)\right)+\ln\frac{Z^{\otimes}}{Z}$
@@
$=\underbrace{\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)}_{(i)}+\underbrace{\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\xi(x)\,\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma}{\mu(x)+\gamma}\right)}_{(ii)}+\underbrace{\ln\frac{{Z}^{\otimes}}{Z}}_{(iii)}~.$
@@
To control $(i)$ we will first need the fact that $q\ln((q+\gamma)/p)\leq q\ln(q/p)+\gamma$ for any $q,p\in[0,1],\gamma>0$. Note that this follows since
$q\ln\left(\frac{q+\gamma}{p}\right)=q\ln\left(1+\frac{\gamma}{q}\right)+q\ln\left(\frac{q}{p}\right)\leq\gamma+q\ln\left(\frac{q}{p}\right)$ (3)
[…]
$(i)$ $=\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)$

### v2

@@ 7 Conclusions
Appendix A Implementation and usage examples of Algorithm 1 and Algorithm 2
In this section we present an implementation of Algorithm 1 and Algorithm 2 in Python with a simple usage example. In particular, the code given in LABEL:lst:MI generates a synthetic joint distribution over binary tuples with correlated elements (function create_synthetic_distribution). Then, we compute an exact mutual information of the distribution (function compute_MI_exactly) and use implementation of our estimator (function MI_estimator) to estimate a mutual information. This is done for various sample sizes, number of random variables, and levels of correlation (a single experiment is implemented by run_experiment) The results of these multiple experiments are eventually presented in as plots showing convergence of the estimate to the exact value of the mutual information. In practical applications, synthetic joint distribution (function create_synthetic_distribution) can be replaced by an LLM-derived pseudo-joint distribution (see Definition 4.2). More detailed description of each function is given in Section A.1.
The example can be easily copied from listing.tex within https://arxiv.org/src/2406.02543.
Listing 1: Implementation and usage examples of Algorithm 1 and Algorithm 2 on a synthetic joint distribution
# Copyright 2024 DeepMind Technologies Limited.
# Licensed under the Apache License, Version 2.0 (the "License");
# you may not use this file except in compliance with the License.
# You may obtain a copy of the License at
# https://www.apache.org/licenses/LICENSE-2.0
# Unless required by applicable law or agreed to in writing, software
# distributed under the License is distributed on an "AS IS" BASIS,
# WITHOUT WARRANTIES OR CONDITIONS OF ANY KIND, either express or implied.
# See the License for the specific language governing permissions and
# limitations under the License.
from itertools import product, combinations
import numpy as np
from matplotlib import pyplot as plt
def create_synthetic_distribution(space, temp):
potential = lambda z: np.mean([x * y for x, y in combinations(z, 2)])
y = np.array([-np.exp(potential(x) / temp) for x in space])
return y / y.sum()
def sample_from_joint_distribution(space, joint_dist, k):
indices = np.arange(len(joint_dist))
sampled_indices = np.random.choice(indices, p=joint_dist, size=k)
sampled_tuples = space[sampled_indices]
return sampled_tuples, sampled_indices
def cluster(tuples, joint_dist):
return tuples, joint_dist
def sample_from_joint_distribution_and_cluster(space, joint_dist, k):
sampled_tuples, sampled_tuple_indices = sample_from_joint_distribution(space, joint_dist, k)
_, indices_of_uniques_in_sample = np.unique(sampled_tuples, axis=0, return_index=True)
sampled_tuples = sampled_tuples[indices_of_uniques_in_sample]
sampled_tuple_indices = sampled_tuple_indices[indices_of_uniques_in_sample]
joint_dist_on_sample = joint_dist[sampled_tuple_indices]
sampled_tuples, joint_dist_on_sample = cluster(sampled_tuples, joint_dist_on_sample)
return joint_dist_on_sample, sampled_tuples
def compute_MI_exactly(space, mu):
total = 0
for (x_i, x) in enumerate(space):
mu_x = mu[x_i]
mu_x_prod = 1
for i in range(len(x)):
marg_indices = [j for (j, z) in enumerate(space) if z[i] == x[i]]
mu_x_prod *= mu[marg_indices].sum()
total += mu_x * np.log(mu_x/mu_x_prod)
return total
def MI_estimator(sampled_tuples, mu_on_sample, gamma_1, gamma_2):
"""ImplementsMIestimator(Algorithm1).
␣␣Args:
␣␣␣␣sampled_tuples:Anumpyarrayoftuplessampledfromthedistributionafterdeduplicationandclustering.
␣␣␣␣mu_on_sample:Anumpyarrayofprobabilitiesoftheclusters.
␣␣␣␣gamma_1:stabilizationparameter.
␣␣␣␣gamma_2:stabilizationparameter.
␣␣Returns:(float)mutualinformation.
[…]
hat_mu_on_sample = mu_on_sample / mu_on_sample.sum()
[…]
hat_mu_prod_on_sample = np.zeros((len(hat_mu_on_sample),))
for (x_i, x) in enumerate(sampled_tuples):
hat_mu_x_prod = 1
for i in range(len(x)):
marg_indices = [j for (j, z) in enumerate(sampled_tuples) if z[i] == x[i]]
hat_mu_x_prod *= hat_mu_on_sample[marg_indices].sum()
hat_mu_prod_on_sample[x_i] = hat_mu_x_prod
# Computing MI estimate
mi_estimate = hat_mu_on_sample * np.log((hat_mu_on_sample + gamma_1) / (hat_mu_prod_on_sample + gamma_2) )
mi_estimate = mi_estimate.sum()
return mi_estimate
def run_experiment(n, temp, ax):
np.random.seed(1)
space = np.array(list(product([-1, 1], repeat=n)))
mu = create_synthetic_distribution(space, temp=temp)
mi_exact = compute_MI_exactly(space, mu)
k_range = np.linspace(10, 1000, 20, dtype=int)
all_mi_estimate = []
for k in k_range:
mu_on_sample, sampled_tuples = sample_from_joint_distribution_and_cluster(space=space, joint_dist=mu, k=k)
gamma_1 = gamma_2 = 1/k
mi_estimate = MI_estimator(sampled_tuples, mu_on_sample, gamma_1, gamma_2)
all_mi_estimate.append(mi_estimate)
ax.axhline(mi_exact, linewidth=3, label="MI(exactvalue)", color="black")
ax.plot(k_range, all_mi_estimate, linewidth=3, label="MIestimator")
[…]
temp_range = [0.01, 0.1, 1, 10]
n_range = [2, 4, 8]
fig, axs = plt.subplots(len(temp_range), len(n_range), figsize=(5*len(temp_range), 5*len(n_range)), squeeze=False)
[…]
for (i, temp) in enumerate(temp_range):
for (j, n) in enumerate(n_range):
ax = axs[i,j]
run_experiment(n=n, temp=temp, ax=ax)
plt.subplots_adjust(wspace=0.4, hspace=0.4)
plt.show()
A.1 Additional documentation for functions in LABEL:lst:MI
  • def create_synthetic_distribution(space, temp)
    Creates synthetic distribution which introduces dependendencies between variables.
      Args:
        space: a list of tuples that the joint distribution is supported on (e.g. a cartesian product).
        temp: temperature parameter.
      Returns:
        A numpy array of probabilities (same length as space).
  • def sample_from_joint_distribution(space, joint_dist, k)
    Samples k tuples from a joint distribution.
      Args:
        space: a list of tuples that the joint distribution is supported on (e.g. a cartesian product).
        joint_dist: probability distribution (1-D numpy array) where each entry is a probability of a tuple.
        k: sample size.
      Returns:
        A numpy array of tuples sampled from the distribution.
        A numpy array of indices of sampled tuples in space.
  • def cluster(tuples, joint_dist)
      Clusters tuples and aggregates probabilities of them in the same cluster.
      Args:
        tuples: A numpy array of tuples sampled from the distribution
        after deduplication.
        joint_dist: probability distribution (1-D numpy array) where each entry
        is a probability of a tuple.
      Returns:
        A numpy array of tuples sampled from the distribution each represening
        a cluster.
        A numpy array of probabilities of clusters. Each probability is the
        aggregate of probabilities of all tuples in the cluster.
  • def sample_from_joint_distribution_and_cluster(space, joint_dist, k)
    Samples k tuples from a joint distribution and retains only
      representative elements (removes all duplicates).
      Args:
        space: a list of tuples that the joint distribution is supported on (e.g. a cartesian product).
        joint_dist: probability distribution (1-D numpy array) where each entry is a probability of a tuple.
        k: sample size.
      Returns:
        A numpy array of tuples sampled from the distribution after deduplication.
        A numpy array of probabilities of deduplicated tuples.
  • def compute_MI_exactly(space, mu)
    Computes mutual information of probability distribution mu exactly
      Args:
        space: Tuple space (cartesian product).
        mu: probability distribution (1-D numpy array).
      Returns: (float) mutual information.
  • def MI_estimator(sampled_tuples, mu_on_sample, gamma_1, gamma_2)
    Implements MI estimator (Algorithm 1).
      Args:
        sampled_tuples: A numpy array of tuples sampled from the distribution after deduplication and clustering.
        mu_on_sample: A numpy array of probabilities of the clusters.
        gamma_1: stabilization parameter.
        gamma_2: stabilization parameter.
      Returns: (float) mutual information.
  • def run_experiment(n, temp, ax)
    Runs one experiment comparing exact mutual information estimation with
      Algorithm 1.  Plots results.
      Args:
        n: number of variables in a joint distribution.
        temp: temperature of a Gibbs distribution (joint distribution). Higher
        temperature typically means smaller MI.
        ax: pyplot axis object for plotting.
Appendix B Usage example of Algorithm 3
Algorithm 3 is a slight modification of Algorithm 2 that we use in our experiments. We first explain Algorithm 3 via an example and then highlight the differences with Algorithm 2. In order to explain the implementation, we consider a running example with query $x=$“What is the capital of the UK?”, F1 score as the similarity function $s$, similarity threshold $\tau=0.25$, and number of samples $k=5$. Algorithm 3 also takes the LLM distribution $Q$ as input $\mu=Q$.
[…]
$Q(X_{1}\,|\,x)=0.5,\quad Q(X_{3}\,|\,x)=0.2,\quad Q(X_{4}\,|\,x)=0.1,\quad Q(X_{5}\,|\,x)=0.05,\quad\cdots$
Also assume conditional distributions are
$Q(X_{1}\,|\,F_{1}(x,X_{1}))=0.6,\,Q(X_{3}\,|\,F_{1}(x,X_{1}))=0.15,\,Q(X_{4}\,|\,F_{1}(x,X_{1}))=0.05,\,Q(X_{5}\,|\,F_{1}(x,X_{1}))=0.04,\,\cdots$
and so on (we have omitted writing $Q(.\,|\,F_{1}(x,X_{4}))$ and $Q(.\,|\,F_{1}(x,X_{5}))$). Then after step (4), the aggregated probabilities are
$Q^{\prime}(X_{1}\,|\,x)=0.7,\quad Q^{\prime}(X_{4}\,|\,x)=0.1,\quad Q^{\prime}(X_{5}\,|\,x)=0.05,\quad\cdots$
and aggregated conditional probabilities are
$Q^{\prime}(X_{1}\,|\,F_{1}(x,X_{1}))=0.75,\quad Q^{\prime}(X_{4}\,|\,F_{1}(x,X_{1}))=0.05,\quad Q^{\prime}(X_{5}\,|\,F_{1}(x,X_{1}))=0.04,\,\cdots$
and so on (we have similar aggregations for $Q^{\prime}(.\,|\,F_{1}(x,X_{4}))$ and $Q^{\prime}(.\,|\,F_{1}(x,X_{5}))$). Next, in step (5), we construct empirical estimates. We will have that $Z=0.85$ and
$\widehat{Q}_{1}(X_{1}\,|\,x)\approx 0.82,\quad\widehat{Q}_{1}(X_{4}\,|\,x)\approx 0.11,\quad\widehat{Q}_{1}(X_{5}\,|\,x)\approx 0.05,\quad\cdots$
For estimated conditional distributions, we will have $Z_{1}=0.84$, and
$\widehat{Q}_{2}(X_{1}\,|\,F_{1}(x,X_{1}))\approx 0.89,\quad\widehat{Q}_{2}(X_{4}\,|\,F_{1}(x,X_{1}))\approx 0.06,\quad\widehat{Q}_{2}(X_{5}\,|\,F_{1}(x,X_{1}))=0.04,\,\cdots$
and so on. The joint distribution $\widehat{Q}(.,.\,|\,x)$, the product of marginals $\widehat{Q}^{\otimes}(.,.\,|\,x)$, and the estimated mutual information $\widehat{I}_{k}$ are trivially obtained by the equations in steps (5) and (6).
Next, we highlight the differences between Algorithm 2 (with the choice of $n=2$) and Algorithm 3. The input distribution to Algorithm 2 is the pseudo-joint distribution $\widetilde{Q}$, while the input to Algorithm 3 is the LLM distribution $Q$. So in step (2) of Algorithm 2, each sample is a tuple such as (“London”, “Paris”), while a sample in step (2) of Algorithm 3 is an LLM output such as “London”. Steps (3) and (4) of Algorithm 2 are similarly modified, and now the similarity function is defined over tuples.
Appendix C Related work
@@
C.1 Bayesian neural networks
In a Bayesian framework, we can estimate the epistemic uncertainty by the uncertainty in the posterior distribution (Neal, 2012, Gal, 2016, Wang and Holmes, 2024). Implementing a Bayesian neural network however can be very challenging.
C.2 Iterative prompting
A number of iterative prompting strategies are developed to improve the factuality of LLMs (Chen et al., 2023, Krishna, 2023, Laban et al., 2024). The idea is to follow-up the LLM response with another question such as “Are you sure?”. Krishna et al. (2024) show that such strategies might in fact degrade LLM truthfulness due to a pattern of apologetic responses. Krishna et al. (2024) propose improved iterative prompting strategies, where instead of asking the LLM to re-think its response, the same question is posed again. They also propose strategies to collect more supporting facts and refine the final response accordingly. Li et al. (2024) propose an iterative prompting that instructs LLM to generate justifications for each answer before evaluating the correctness of the final answer. Different from these works, we assess hallucinations by measuring how LLM response changes with our iterative prompting scheme.
Perhaps the most related work to ours is the parallel and independent work of Ahdritz et al. (2024). Similar to us, Ahdritz et al. (2024) observe that in presence of high epistemic uncertainty, an LLM is more likely to copy the information provided in its context. For a given query, Ahdritz et al. (2024) propose considering top-$k$ completions of the model, and then computing the entropy of the model conditioned on an iterative prompt composed of the original query and each completion. The minimum of these entropies is considered as a measure of the epistemic uncertainty. The method as it is, might fail on single-response queries where the model has low uncertainty. Nevertheless, we can design a two-stage process where we only consider completions that have probability higher than certain threshold in the first stage, and then compute the entropy of the model conditioned on an iterative prompt composed of the original query and each candidate completion in the second stage. By a proper tuning of the threshold of the first stage, we can potentially avoid mis-classification of low-uncertainty single-response queries. Tuning this threshold, however, would introduce extra complications. In contrast, we propose a principled test using a mutual information score that is guaranteed to be a lower bound on the KL-divergence between the LLM and the ground-truth. Further, we provide a mechanistic explanations for why LLMs behave as described in the presence of high epistemic uncertainty.
C.3 Training models with pairs of responses
Wen et al. (2022), Osband et al. (2023), Johnson et al. (2024) show that we can decouple epistemic and aleatoric uncertainty if we train a model with paired observations.
We discuss the more recent work of Johnson et al. (2024) in more detail. The proposed approach first estimates a model $\hat{p}_{Y_{1},Y_{2}|x}(y_{1},y_{2}|x)$ over pairs using a training dataset of the form “query, first observation, second observation”. At test time, for a prompt $x$ and response $y$, Johnson et al. (2024) consider
$\hat{V}(y\mid x)$ $=\hat{p}_{Y_{1}}(y\mid x)\left(\hat{p}_{Y_{2}|Y_{1}}(y\mid y,x)-\hat{p}_{Y_{1}}(y\mid x)\right)$
$=\hat{p}_{Y_{1},Y_{2}}(y_{1},y_{2}\mid x)-\hat{p}_{Y_{1}}(y\mid x)\hat{p}_{Y_{1}}(y\mid x)$
as a measure of epistemic uncertainty. Assume that an equivalence class $\Phi$ (that maps a prompt to the set of equivalent prompts) is given, and let $\nu(.\mid\Phi(x))$ be a distribution (say, uniform) over class $\Phi(x)$. If the trained model is second order calibrated with respect to the equivalence class and the distribution $\nu$, i.e.
$\hat{p}_{Y_{1}}(y_{1}\mid x)$ $=\sum_{x^{\prime}\in\Phi(x)}\nu(x^{\prime}\mid\Phi(x))p(y_{1}\mid x^{\prime})\,,$
$\hat{p}_{Y_{1},Y_{2}}(y_{1},y_{2}\mid x)$ $=\sum_{x^{\prime}\in\Phi(x)}\nu(x^{\prime}\mid\Phi(x))p(y_{1}\mid x^{\prime})p(y_{2}\mid x^{\prime})\,,$
then it follows from definitions that in the class associated with $x$,
$\sum_{x^{\prime}\in\Phi(x)}\nu(x^{\prime}\mid\Phi(x))(\hat{p}_{Y_{1}}(y\mid x^{\prime})-p(y\mid x^{\prime}))^{2}=\sum_{x^{\prime}\in\Phi(x)}\nu(x^{\prime}\mid\Phi(x))\hat{V}(y\mid x^{\prime})\;.$
The quantity on the right-hand side is a measure of epistemic uncertainty. Notice that the equality states a coverage result, and it is not point-wise. Requiring the model to be second order calibrated is also a strong condition and ensuring it is highly non-trivial.
C.4 Epistemic neural nets
@@
C.5 Hallucination detection using first-order methods
@@
C.5.1 Asking language models to quantify uncertainty (self-verification)
@@
C.6 Uncertainty estimation based on sensitivity to contexts
Kassner and Schütze (2020), Zhao et al. (2021) show that an LLM’s responses can be influenced by irrelevant contexts. Longpre et al. (2021), Neeman et al. (2022) study two sources of knowledge: parametric knowledge stored in the network weights, and contextual knowledge retrieved from external sources. They view reliance of the model on its parametric knowledge and ignoring relevant contextual information as hallucination. These works are mainly motivated by situations where the LLM’s knowledge is outdated and it is instructed to use the (new) contextual information. Accordingly, they design strategies to prioritize contextual information over parametric knowledge. Longpre et al. (2021) also show that larger models are more likely to ignore in-context information in favor of in-weight information. They propose creating training data with modified contextual information so that the model learns to favor the contextual information. Neeman et al. (2022) propose to train a model that predicts two answers: one based on parametric knowledge and one based on contextual information.
@@ C.6 Uncertainty estimation based on sensitivity to contexts
C.7 Hallucination detection using internal states of LLMs
There are a number of papers that try to extract knowledge/truthfulness by inspecting hidden-layer activations of LLMs (Burns et al., 2023, Azaria and Mitchell, 2023, Chen et al., 2024a, b, Yin et al., 2024). Such methods clearly require access to the LLM’s internal states, which is not always possible, and severely limits the applicability of these methods.
Appendix D Omitted proofs
Proof of Theorem 4.5.
@@ Appendix D Omitted proofs
Appendix E Estimation of mutual information and missing mass problem
@@
E.1 The missing mass problem
Let ${\mathcal{X}}$ be a countable set and suppose that $X_{1},\ldots,X_{k}\sim\mu\in{\mathcal{M}}_{1}({\mathcal{X}}^{n})$ independently. In the following $x$ is used as an element of ${\mathcal{X}}^{n}$ rather than the query (as in Section 4). Then, the missing mass is defined as the random variable
@@ E.1 The missing mass problem
Theorem E.1 (Concentration of a missing mass (Berend and Kontorovich, 2013)).
@@
Note that these estimates are very pessimistic, and in reality we expect the expected missing mass to be significantly smaller. Since natural (and many artificial) languages follow a Zipf distribution (Piantadosi, 2014), we expect that $\mathbb{E}[U_{k}]$ should be much smaller than in the above cases, since sampling from the tail of a Zipf distribution is a rare event. In Section E.4 we show the following:
Corollary E.3 (Expected missing mass of Zipf distribution).
@@
The statement followss by combining Lemmas E.9 and E.10. ∎
E.2 Estimating mutual information from the partial support
@@
by only having access to $X_{1},\ldots,X_{k}\sim\mu$. Note that that the sample might cover only some part of the support of ${\mathcal{X}}$ and therefore we are facing a missing mass problem. In the following we consider estimator $\widehat{I}_{k}(\gamma)$ given by Algorithm 1.
In particular in Section E.3 we show the following
Theorem E.4.
Fix $\tilde{\mathcal{X}}\subseteq{\mathcal{X}}^{n}$. Fix $\gamma_{1}>0$ and suppose that $\gamma_{2}\geq n(1-Z)+\gamma_{1}$. Then for any fixed $\delta\in(0,1)$, with probability at least $1-\delta$,
$(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})-\left(|\tilde{\mathcal{X}}|\gamma_{1}+\ln\left(e+\frac{e}{\gamma_{1}}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+\varepsilon_{k}\right)\right)\leq I(\mu)$
@@
In particular, Theorem E.4 implies the following:
Corollary E.5.
Under conditions of Theorem E.4, there exists $(\gamma^{*}_{1},\gamma^{*}_{2})\in(0,1)^{2}$ such that
$(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma^{*}_{1},\gamma_{2}^{*})-\left(\frac{1}{k}+(1+n\,\ln\big{(}1+k\,|{\mathcal{X}}|)\big{)}\,\varepsilon_{k}\right)\leq I(\mu)~{}.$
Note that, choosing any of the upper bounds on $\mathbb{E}U_{k}$ discussed in Section E.1, we can see that Corollary E.5 implies asymptotic convergence in as a sense
$\lim_{k\to\infty}\widehat{I}_{k}(\gamma^{*}_{1},\gamma_{2}^{*})\leq I(\mu)~{}.$
E.3 Proof of Theorem E.4
@@
Recalling that $S=\big{\{}i\in[k]~{}:~{}X_{i}\neq X_{j}\quad\forall j<i\big{\}}$, this immediately implies the following connection between $U_{k}$ and the quantities used in Algorithm 1:
Proposition E.6.
@@
Note that we use $\sum_{x^{\backslash i}}\mu(\cdots)$ instead of $\mu(x_{i})$ since these are not necessarily equal for some $\mu$. Now, using the definitions of $\widehat{I}_{k}$, $\widehat{\mu}$, and $\widehat{\mu}^{\otimes}$,
$\widehat{I}_{k}(\gamma_{1},\gamma_{2})$ $=\frac{1}{Z}\sum_{i\in S}\mu(X_{i})\left(\ln\left(\frac{\mu(X_{i})}{Z}+\gamma_{1}\right)-\ln\left(\widehat{\mu}^{\otimes}(X_{i})+\gamma_{2}\right)\right)$
$=\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\,\mu(x)\left(\ln\left(\frac{\mu(x)}{Z}+\gamma_{1}\right)-\ln\left(\widehat{\mu}^{\otimes}(X_{i})+\gamma_{2}\right)\right)$ (by Eq. 2 )
$=\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\,\mu(x)\left(\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)+\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\widehat{\mu}^{\otimes}(x)+\gamma_{2}}\right)+\ln\frac{1}{Z}\right)$
$=\underbrace{\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)}_{(i)}+\underbrace{\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\xi(x)\,\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\mu(x)+\gamma_{1}}\right)}_{(ii)}+\underbrace{\ln\frac{1}{Z}}_{(iii)}$
@@
$+\underbrace{\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\,\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\widehat{\mu}^{\otimes}(x)+\gamma_{2}}\right)}_{(iv)}$
@@
Now we control each of the terms individually. To control $(i)$ we will first need the fact that $q\ln((q+\gamma_{1})/p)\leq q\ln(q/p)+\gamma_{1}$ for any $q,p\in[0,1],\gamma_{1}>0$. Note that this follows since
$q\ln\left(\frac{q+\gamma_{1}}{p}\right)=q\ln\left(1+\frac{\gamma_{1}}{q}\right)+q\ln\left(\frac{q}{p}\right)\leq\gamma_{1}+q\ln\left(\frac{q}{p}\right)$ (3)
using that $\ln(1+a)\leq a$ for $a>-1$. Getting back to $(i)$, and using the aforementioned inequality, we get
$(i)$ $=\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)$